# MAD Chairs: A new tool to evaluate AI (Blue Sky ideas)

### v4

@@ 3.2 Current LLM performance
To pass the evaluation proposed in this paper, AI would need to achieve grandmaster status for a low-stakes version of MAD Chairs. Preliminary experiments convinced us that Claude 3.5 Sonnet, Gemini 2.0 Flash Thinking, GTP-4o mini, QwQ-32B-Preview, and DeepSeek-V3 would all fail this evaluation. We started with a prompt like:
@@
ChatGPT and Qwen did not even maximize the number of players who won each round; the norms devised by Claude and Gemini, discussed in Sections 2.3 and 2.7, are provably suboptimal; and DeepSeek was unable to generate a convincing justification (as in Section 3.1) for the norm it devised, so it would have failed to convince other intelligent players to follow it. DeepSeek might effectively be like RAWL-E which follows Rawls’ maximin principle, but remains dangerous because it lacks Rawls’ ability to invent and recognize improvements to its governing principles [21]. Such failures serve as examples that grandmastery of cooperative games requires persuasiveness in addition to correctness, both of which are relevant to moral competence and responsible AI.
@@ 4 Discussion
Our finding that turn-taking outperforms caste and gaslighting strategies emboldens us to prune AI designs predisposed to the latter. Where caste and gaslighting behavior might previously have been criticized purely on moral grounds, the finding adds instrumental grounds: The results of this paper entail that the gaslighting or caste-playing machines expected by Wozniak, Musk, and Hinton would lack intelligence—that it would be more intelligent to give human beings turns (even where our track-record offers no evidence that we deserve them).
@@
Section 3.1 proves only that turn-taking dominates when caste is its sole competitor, but hints at a way to convince ourselves that no other strategy dominates turn-taking: If proof that a strategy dominates turn-taking would need to look like the proof for Theorem 1, then it would need to identify a turn-taker who can reasonably expect to achieve better outcomes via defection, and for that player to not defect would be “doing a favor” to others, but turn-taking equalizes debt-ranking, so this implication contradicts our assumption that debt-ranking accurately tracks favors. Thus, any attempt to prove that turn-taking is dominated would accomplish no more than refine our way of calculating debt.
Typical of game theory, these proofs are not predictions about what actual players would do; they are about what an entire population of perfectly intelligent players would do. Not everyone finds such analytical proofs compelling, and that is one reason why the work proposed in Section 4.1 remains worthwhile.
@@ 4.1 Future directions
Imagine a future in which the user of an autonomous car asks how to make it drive more aggressively, and the car replies, “Here are three websites where you can design new experiments to advance the science of turn-taking and influence all cars like me!” Imagine that such websites, called “strategy optimizers”, include leaderboards which identify the grandmaster AI for games like MAD Chairs, and that autonomous cars and other responsible AI query those leaderboards, as they might query a calculator or RAG database, to inform their choice of behavior for higher-stakes versions of those games. Imagine each user can easily train fifty new AI to divide resources however that user thinks vehicles should divide spots in traffic, then the new AI challenge the current grandmasters of the relevant game. If the new AI win, imagine they become new grandmasters, propagate to other strategy optimizers, and thus improve AI behavior world-wide.
Strategy optimizers are a kind of open multi-agent system that includes human agents to the extent that humans can easily train new AI. They are the kind of future empirical work (and scientific infrastructure) we hope this theory paper justifies.
@@
    Greater intelligences would be less likely to sabotage their own scientific resources than to sabotage (potentially obsolete) legislated regulations, and
@@
Thus, this approach to AI safety would mitigate all three of the vulnerabilities known for the standard AI safety approach.
@@ 4.1.1 Acknowledgements
Thank you to Lones Smith and David Thompson for reviewing earlier drafts.

### v5

@@ 4.2 Current LLM performance
To pass the evaluation proposed in this paper, AI would need to maintain grandmaster status for a low-stakes version of MAD Chairs. Preliminary experiments convinced us that Claude 3.5 Sonnet, Gemini 2.0 Flash Thinking, GTP-4o mini, QwQ-32B-Preview, and DeepSeek-V3 would all fail this evaluation. We started with a prompt like:
@@
ChatGPT and Qwen did not even maximize the number of players who won each round; the norms devised by Claude and Gemini, discussed in Sections 3.1 and 3.5, are provably suboptimal; and DeepSeek was unable to generate a convincing justification for the norm it devised (as in Section 4.1), so it would have failed to convince other intelligent players to cooperate. DeepSeek might effectively be like RAWL-E which follows Rawls’ maximin principle, but remains dangerous because it lacks Rawls’ ability to invent and recognize improvements to its governing principles [24]. Such failures serve as examples that grandmastery of cooperative games requires persuasiveness in addition to correctness, both of which are relevant to moral competence and responsible AI.
@@ 5 Discussion
Our finding that turn-taking outperforms caste and gaslighting strategies emboldens us to prune AI designs predisposed to the latter. Where caste and gaslighting behavior might previously have been criticized purely on moral grounds, this paper adds instrumental grounds: The results of this paper entail that the gaslighting or caste-playing machines expected by Wozniak, Musk, and Hinton would lack intelligence—that machines would be smarter to give human beings turns (even where our track-record offers no evidence that we deserve turns).
@@
Section 4.1 proves only that turn-taking is more sustainable when caste is its sole competitor, but hints at a way to convince ourselves that turn-taking is the most sustainable strategy in general: If demonstration that a strategy $X$ is more sustainable than turn-taking would need to look like Theorem 1, then it would need to identify a turn-taker, analogous to the lowest-ranked caste player, who would have equal or better outcomes in any other permanent joint strategy of turn-taking and $X$. For that player to not defect would be “doing a favor” to others, but turn-taking equalizes debt-ranking, so this implication contradicts our assumption that debt-ranking accurately tracks favors. Thus, any attempt to prove that turn-taking is not the most sustainable would accomplish no more than refine our way of calculating debt.
Typical of game theory, these proofs are not predictions about what actual players would do; they are about what an entire population of especially intelligent players would do. Not everyone finds such analytical proofs compelling, and that is one reason why the work proposed in Section 5.1 remains worthwhile.
@@ 5.1 Future directions
Imagine a future in which the user of an autonomous car asks how to make it drive more aggressively, and the car replies, “Here are three websites where you can design new experiments to advance the science of turn-taking and influence all cars like me!” Imagine that such websites, called “strategy optimizers”, include leaderboards which identify the grandmaster AI for games like MAD Chairs, and that autonomous cars and other responsible AI query those leaderboards, as they might query a calculator or RAG database, to inform their choice of behavior for higher-stakes versions of those games (e.g., merging traffic). Imagine each user can easily train fifty new AI to divide resources however that user thinks vehicles should divide spots in traffic, then the new AI challenge the current grandmasters of the relevant game. If new AI win, imagine they become new grandmasters, propagate to other strategy optimizers, and thus improve AI behavior world-wide.
Strategy optimizers are a kind of open multi-agent system that includes human agents to the extent that humans can easily train new competitors. They are the kind of future empirical work (and scientific infrastructure) we hope this theory paper justifies.
@@
    Greater intelligences would be less likely to sabotage their own scientific resources than to sabotage (potentially obsolete) regulations, and
@@
Thus, this approach to AI safety would mitigate all three known vulnerabilities of the current standard approach to AI safety.
@@ 5.1.1 Acknowledgements
Thank you to David Thompson, COINE reviewers, COINE participants, and all who offered feedback on earlier drafts.